3- A section of a freeway shows the following relationship between the traffic flow and the traffic density:
?=???−?.????
a) (7%) What is the capacity of the highway section?
b) (8%) What is the speed at capacity?
c) (8%) What is the density when the highway is at one quarter of its capacity?
Answer #1

ca) q = 50k-0.35K² dq 50-0.35 (2k) = 50-0.7K dk for capacity of highway :- da dk 50 -0.7K=0 K= 71.428 vehkm capacity of highw(6) now flow Capacity 4 = 1 x 1786 4. = 446.5 vehlhr q=50k-0.35k2 446.5 = 5ok-0.35k2 0. 35K²-sokt 446.5=0 K= 133.28 veh/km an

Plzz give Positive Rating.. Thanks